Let F be the field of complex numbers and let T be the function from F

3

into F

3 defined by


T(x1, x2, x3) = (x1 − x2 + 2x3 ,2x1 + x2 , − x1 − 2 x2 + 2x3 )

i. Verify that T is a linear transformation.

ii. If (a, b, c) is a vector in F

3, what are the conditions on a, b and c that the vector bein the range of T? What is the rank of T?

iii. What are the conditions on a, b and c that the vector (a, b, c) be in the null space of T? What is the nullity of T?


1
Expert's answer
2021-01-04T20:16:10-0500

"T:F_3\rightarrow F_3\\
(x_1,x_2,x_3)\rightarrow (x_1 − x_2 + 2x_3 ,2x_1 + x_2 , − x_1 − 2 x_2 + 2x_3 )"

"=\begin{bmatrix}1&-1&2\\2&1&0\\-1&-2&2\end{bmatrix}\times\begin{bmatrix}x_1\\x_2\\x_3\end{bmatrix}=AX \;[let]"

where "A=\begin{bmatrix}1&-1&2\\2&1&0\\-1&-2&2\end{bmatrix},X=\begin{bmatrix}x_1\\x_2\\x_3\end{bmatrix}"

i.e. "T(X)=AX"


(i)

Let "X,Y\in F_3, c\in \mathbb{C}"

"\therefore T(aX+Y)=A(aX+Y)=aAX+AY=aT(X)+T(Y)"

Hence "T" is a linear transformation.


(ii)

Claim:- "\{A_1=\begin{bmatrix}1\\2\\-1\end{bmatrix},A_2=\begin{bmatrix}-1\\1\\-2\end{bmatrix}\}" is a Basis of "Im(T)" .

Proof:-

Any element of "Im(T)" can be expressed by columns of "A \;\;i.e. \;A_1, A_2,A_3"

"\therefore Im(T)=span(\{A_1, A_2,A_3\})"

"A_3=\frac{2}{3}A_1-\frac{4}{3}A_2"

"\therefore A_3\in span(\{A_1,A_2\})"

"\Rightarrow span(\{A_1,A_2\})= span(\{A_1,A_2,A_3\})=Im(T)" ...(1)

Now "A_1,A_2" are independent [ as "\exists no\;c\in \mathbb{C}\ni A_1=cA_2" ] ...(2)


So, from above two facts, we can conclude that ,

"\{\begin{bmatrix}1\\2\\-1\end{bmatrix},\begin{bmatrix}-1\\1\\-2\end{bmatrix}\}" is a Basis of "Im(T)" ...(Proved)


Therefore "\pmb{rank(T)=dim(Im(T))=2}"


(iii)

Null Space of "\;T=Null(T)=\{X\in F_3 | AX=0\}"

So,

"\;\;\;\;\;AX=0\\
\Rightarrow \begin{bmatrix}1&-1&2\\2&1&0\\-1&-2&2\end{bmatrix}\begin{bmatrix}x_1\\x_2\\x_3\end{bmatrix}=0\\
\Rightarrow \begin{bmatrix}1&0&2/3\\0&1&-4/3\\0&0&0\end{bmatrix}\begin{bmatrix}x_1\\x_2\\x_3\end{bmatrix}=0\\
\Rightarrow x_1+\frac{2x_3}{3}=0, \;x_2-\frac{4x_3}{3}=0"

[let "x_3=k"]

"\Rightarrow X\in\{k\begin{bmatrix}-2/3\\4/3\\1\end{bmatrix}: k\in \mathbb{C}\}"

"\therefore Null(T)=\{k\begin{bmatrix}-2/3\\4/3\\1\end{bmatrix}: k\in \mathbb{C}\}"

Clearly Null(T) is generated by 1 element.

Hence, "\pmb{Nullity\;\;of \;\;T=dim(Null(T))=1}".